Circular No. 8277

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004P IN UGC 8561 Further to IAUC 8274, J. Burket and W. Li report the LOSS discovery,

on a KAIT image taken on Jan. 23.5 UT (mag ∼ 16.3), of an apparent supernova located at α = 13^{h}34^{m}57^{s}.76, δ = +34^{o}02′33′′.7 (equinox 2000.0), which is 6′′.0 east and 4′′.5 south of the nucleus of UGC 8561. The new object was confirmed at mag ∼ 16.5 on an image taken on Jan. 27.5 by M. Schwartz and P. Holvorcem with the 0.8-m Tenagra II telescope. A KAIT image taken on 2003 June 17.2 showed nothing at this position (limiting mag ∼ 19.0).

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004P, obtained by P. Berlind on Jan. 27.53 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova ∼ three weeks past maximum.

SUPERNOVAE 2004J, 2004K, AND 2004L Matheson et al. also report that a spectrum of SN 2004J (cf. IAUC

8272), obtained as above on Jan. 27.20 UT, shows it to be a type-Ia supernova about two weeks past maximum. Narrow emisson lines superimposed on the spectrum of the supernova indicate a recession velocity for the host galaxy of 7570 km/s. A spectrogram of SN 2004K (cf. IAUC 8273), obtained by Berlind on Jan. 27.55, shows it to be a type-Ia supernova about a week past maximum; adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 10832 km/s for the host galaxy, the supernova expansion velocity is 11800 km/s for Si II (rest 635.5 nm). The spectralfeature age (Riess et al. 1997, A.J. 114, 722) of SN 2004K is 8 ± 2 days after maximum light. A spectrogram of SN 2004L (cf. IAUC 8274), obtained on Jan. 27.40, shows it to be a type-Ia supernova near maximum; adopting the NED recession velocity of 9686 km/s for the host galaxy, the supernova expansion velocity is 10000 km/s for Si II (rest 635.5 nm). The spectral-feature age of SN 2004L is 1 ± 2 days before maximum light.

C. L. Gerardy reports that a low-resolution optical spectrogram (obtained as reported on IAUC 8275 for SN 2004I) of SN 2004L, obtained on Jan. 25.29 UT, shows it to be a type-Ia supernova shortly after maxiumum light; adopting from NED a redshift of 9686 km/s for MCG +03-27-38, the expansion velocity of the Si II feature (rest 635.5 nm) is 9600 km/s.

2004 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green